# Graph Neural Networks for Jamming Source Localization

### v1

@@ Abstract
Graph-based learning has emerged as a transformative approach for modeling complex relationships across diverse domains, yet its potential in wireless security remains largely unexplored. In this work, we introduce the first application of graph-based learning for jamming source localization, addressing the imminent threat of jamming attacks in wireless networks. Unlike geometric optimization techniques that struggle under environmental uncertainties and dense interference, we reformulate the localization as an inductive graph regression task. Our approach integrates structured node representations that encode local and global signal aggregation, ensuring spatial coherence and adaptive signal fusion. To enhance robustness, we incorporate an attention-based Graph Neural Network (GNN) that adaptively refines neighborhood influence and introduces a confidence-guided estimation mechanism that dynamically balances learned predictions with domain-informed priors. We evaluate our approach under complex Radio Frequency (RF) environments with varying sampling densities and signal propagation conditions, conducting comprehensive ablation studies on graph construction, feature selection, and pooling strategies. Results demonstrate that our novel graph-based learning framework significantly outperforms established localization baselines, particularly in challenging scenarios with sparse and obfuscated signal information. Our code is available at https://github.com/daniaherzalla/gnn-jamming-source-localization.
@@ 1 Introduction
Graphs serve as a fundamental framework for representing complex relationships and interactions in real-world systems. Their versatility allows for the sophisticated modeling of diverse data types, driving their adoption across scientific and industrial domains such as molecular property prediction, social recommendation systems, and traffic forecasting. Many of the most successful machine learning applications, such as image, text, and speech processing, can be viewed as specialized cases of graph-based learning, underscoring the broad relevance and transformative potential of graph representations in advancing machine learning methodologies [23]. Although graph-based learning has been explored in numerous domains, its application to wireless security remains relatively underexplored. In particular, jamming source localization, a task essential for mitigating the imminent threat posed by disruptive sources of interference to the availability of wireless communication, is yet to be explored through this application, presenting a novel avenue for research and application. In this study, we investigate the application of a novel graph-based learning framework to perform an inductive graph-level regression task to predict the location of a jamming source in complex RF operating environments.
@@
The ubiquitous and growing dependence on wireless networks for applications ranging from everyday connectivity to mission-critical operations introduces significant security vulnerabilities. As jamming attacks transmit intentional interference signals across communication frequencies, they effectively lead to the degradation and severing of wireless links [15]. The repercussions of such attacks are severe, leading to failures in mission coordination, disruptions in essential services, and even operational hazards [28]. While countermeasures such as frequency hopping have been employed, their effectiveness is limited against advanced jammers such as reactive, follow-on, and barrage that adaptively pursue target frequencies and create dense interference across multiple channels [16]. Unlike avoidance techniques that passively adjust the network to evade interference, jamming source localization can provide a direct mitigation strategy. By accurately identifying the jammer’s position, network administrators can deploy targeted countermeasures such as physical neutralization, geofencing, or jamming suppression, restoring network reliability regardless of the changing dynamics of the jammer’s strategy [28].
@@
Classical jamming localization methods rely on geometric and optimization-based techniques to estimate jammer positions. However, they degrade in real-world RF environments due to noise and multipath effects [13]. Their dependence on idealized propagation models limits adaptability to stochastic RF dynamics. To overcome these limitations, we introduce a novel graph-based formulation of the jamming localization problem, leveraging attention-based GNNs to adaptively extract spatial and signal-related patterns from measurements. Additionally, we propose Confidence-guided Adaptive Global Estimation (CAGE), a confidence-guided estimation mechanism that dynamically balances GNN-based predictions with domain-informed priors, improving robustness across varying deployment conditions. Rather than treating localization as a geometric optimization problem, we redefine it as a graph regression task, where node features encode key RF and spatial characteristics, and the graph structure captures local and global signal dependencies. Our contributions are as follows:
@@
  • We present the first application of GNNs to jamming source localization in wireless networks by reformulating the problem as a graph regression task.
@@ 2.1 Jamming Source Localization
Jamming source localization has been widely studied under range-free and range-based algorithms. Range-free algorithms determine location information using network topology-related properties, without relying on the physical characteristics of incoming signals [10]. These methods are useful in infrastructure-limited environments, however, they suffer from degraded accuracy in sparse or unevenly distributed networks [28, 25]. In contrast, range-based methods utilize geometric optimization to estimate the distance to the source using the measurement of various physical properties such as Received Signal Strength Indicator (RSSI), Time of Arrival (ToA), and Angle-of-Arrival (AoA[10]. While these methods generally achieve higher accuracy, obtaining measurements such as ToA and AoA rely on specialized and calibrated hardware and are susceptible to errors where multipath effects introduce significant biases [29].
Many of the existing jammer localization approaches are validated under theoretical propagation models, such as free-space path loss, which fails to capture the complexities of real-world settings. Additionally, many evaluations do not encompass a diverse array of network attack scenarios, including irregular node placements, sparse samples, and long-range jamming attacks [14]. To address these limitations, we evaluate our method on diverse network configurations, leveraging the Log Distance Path Loss (LDPL[6] model to account for realistic signal propagation conditions (see Appendix 0.F).
As our method leverages spatial and signal information exclusively, we focus on methods that can similarly harness this information for localization. We benchmark our approach against established range-free and range-based localization techniques: Weighted Centroid Localization (WCL[25], Least Squares (LSQ[28], Path Loss (PL[12], Maximum Likelihood Estimation (MLE[12], and Multilateration (MLAT[30]. These baselines serve as reference points to assess the robustness and adaptability of our graph-based framework.
@@ 2.2 Graph Neural Networks for Localization
GNNs have demonstrated strong capabilities in regression tasks by capturing spatial dependencies within graph-structured data, such as in molecular property prediction and material science [11]. Beyond these domains, GNNs have also been applied to pose regression problems, including camera pose estimation [22] and human pose tracking [31], where they refine node and edge representations to improve motion prediction and spatial consistency. Their effectiveness in localization tasks has also gained attention, particularly in wireless and sensor network applications. GNNs have been applied to RF-based localization, including WiFi fingerprinting-based indoor localization [8], where GNNs leverage RSSI signals to construct graphs and improve positioning accuracy. More recent works extend GNNs for network localization beyond Wi-Fi, addressing challenges such as dynamic network topologies and mobility-induced signal variation [6].
While these advancements highlight the growing potential of GNNs for network localization, their application to jamming source localization remains unexplored. This gap presents an opportunity to adapt GNNs to interference localization, addressing unique challenges in adversarial conditions.
@@ 3 Problem Definition
The problem addressed in this study is the localization of a wireless jammer, an adversarial interference source that disrupts communications by emitting high-power signals to degrade legitimate wireless communication.
Network Configuration
@@
where $s_{i}^{t}$ denotes a measurement taken at position $\mathbf{x}_{i}^{t}$ with an associated noise floor value $\eta_{i}^{t}$ in dBm at time $t$. Note that $\eta_{i}^{t}$, referred to as the noise floor, is commonly treated as jamming signal strength, where here it represents the combined effect of jammer interference and baseline environmental noise floor. Signal attenuation and propagation effects are modeled using the empirical LDPL model to simulate Non-line-of-sight (NLOS) conditions; refer to Appendix 0.F for details.
@@
Jammer Characteristics
The jammer is located at an unknown position $\mathbf{x}_{j}\in\mathcal{A}$ and emits interference signals that elevate the noise floor of nearby devices. The interference strength varies spatially due to distance attenuation and propagation effects such as shadowing, reflections, and multipath fading. The sampled region, denoted as $\mathcal{R}\subseteq\mathcal{A}$, is the subset of $\mathcal{A}$ where devices collect noise floor measurements. Although the jammer’s position satisfies $\mathbf{x}_{j}\in\mathcal{A}$, it does not necessarily hold that $\mathbf{x}_{j}\in\mathcal{R}$, i.e., $\mathbf{x}_{j}\in\mathcal{A}\setminus\mathcal{R}$ is possible. We evaluate localization methods in scenarios where the jammer affects areas beyond the sampled region, testing their ability to infer positions outside direct measurement zones.
Objective
Given the set of samples collected by $N$ devices, denoted as $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\dots,\mathcal{S}_{N}\}$, the goal is to infer $\mathbf{x}_{j}$ based on the spatial distribution of measured noise floor levels. This problem is inherently challenging due to incomplete and noisy measurements, environmental uncertainty affecting signal propagation, complex spatial correlations between noise floor levels and jammer interference, and the need to generalize beyond observed regions where no direct measurements are available.
@@ 4 Graph-Structured Learning for Jammer Localization
We structure the above problem as a graph representation to systematically model spatial relationships and signal attenuation. Given the set of collected measurements $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\dots,\mathcal{S}_{N}\}$, we represent the sampled signal space as a graph $G=(V,E)$, where nodes in $V$ correspond to individual measurement instances $s_{i}^{t}\in\mathcal{S}_{i}$, and edges $E$ define spatial relationships between them. Edges are established using K-Nearest Neighbors (KNN), where for each node $v_{i}$, an edge set is constructed as $E=\{(v_{i},v_{j})\mid v_{j}\in\mathcal{N}_{k}(v_{i})\}$, where $\mathcal{N}_{k}(v_{i})$ denotes the set of $k$ nearest neighbors of $v_{i}$ in the Euclidean space. To enforce the spatial attenuation principle of jamming signals, each edge $(v_{i},v_{j})$ is assigned a weight $w_{ij}$ that decays exponentially with the Euclidean distance as:
@@ 4.1 Learning Spatial Relations with Attention-Based Graphs
Graph-based learning enables the model to capture structured dependencies in the jammer interference field, allowing for improved generalization across spatial regions. However, effective learning in this setting requires handling two key challenges: (1) signal measurements are inherently noisy due to environmental conditions, leading to unreliable observations, and (2) the importance of neighboring nodes varies depending on both their spatial proximity and their reliability in capturing interference effects. Traditional GNN architectures, such as Graph Convolutional Networks [7], rely on uniform neighborhood aggregation and struggle to capture the varying reliability of spatially adjacent nodes. Our approach follows Graph Attention Network (GAT[24], incorporating an adaptive weighting mechanism that dynamically refines neighborhood influence, ensuring that message passing prioritizes nodes carrying more reliable signal information while attenuating contributions from potentially misleading observations.
@@ 4.2 Supernode-Guided Adaptive Estimation
Global Context Encoding with Domain-Guided Priors
@@
Figure 1: Overview of the proposed jammer localization framework. A graph is constructed from noise floor measurements, where nodes represent signal instances, and edges capture spatial relationships. The encoder processes the graph to learn spatial correlations. The final jammer position is estimated through an adaptive combination of the GNN-based prediction and a WCL prior, controlled by a learned confidence weight $\alpha$.
Confidence-Guided Adaptive Position Estimation
@@
where $\boldsymbol{\alpha}\in\mathbb{R}^{5}$ is a five-dimensional confidence vector, with each $\alpha_{d}$ corresponding to one of the five output components in the normalized angular representation. The parameters $\mathbf{W}_{\alpha}\in\mathbb{R}^{5\times F^{\prime}}$ and $\mathbf{b}_{\alpha}\in\mathbb{R}^{5}$ are learnable, and sigmoid $\sigma(\cdot)$ ensures that $10<\alpha_{d}<$ 1. Finally, the final predicted jammer position is computed as:
@@ 4.3 Training Strategy and Loss Function
Encouraging Independent Learning of the GNN Regressor
@@
Note that, in experimental evaluation, we refer to our proposed method as CAGE. For clarity in comparisons, Multilayer Perceptron (MLP), GCN, Principal Neighbourhood Aggregation (PNA) and GAT operate on graph $G$ and are trained Equation (15) with the final estimate given by $\hat{\mathbf{x}}_{\text{GNN}}$, while CAGE is evaluated on the augmented graph $G^{\prime}$ and the adaptive confidence-weighted estimation, trained with Equation (14).
@@ 5 Experimental Evaluation
We evaluate CAGE across static and dynamic environments under the LDPL model. In the static setting (Section 5.1), fixed nodes with varying positions and densities are considered, with the jammer positioned either inside or outside the sampled region $\mathcal{R}$ to study spatial effects. In the dynamic setting (Section 5.2), a device moves in 3D space, approaching and encircling the jammer to assess localization accuracy across angles and distances. Ablation studies in Section 5.3 and Appendix 0.A evaluate the impact of node features, edge definitions, graph augmentations, global pooling mechanisms, and downsampling techniques. Further details on data generation, spatial arrangements, trajectory modeling, and signal propagation are provided in Appendix 0.F, and additional analysis of confidence weighting is presented in Appendix 0.B.
We compare CAGE against classical methods (WCL [28], LSQ [28], PL [12], MLAT [30], MLE [12], MLP) and graph learning-based methods (GCN [7], PNA [3], GAT [24]). As previously described, while MLP, GCN, PNA, and GAT operate on $G$, CAGE leverages augmented graph $G^{\prime}$ with the supernode, incorporating graph attention mechanisms alongside confidence-weighted estimation for adaptive localization. Models are trained using AdamW with a one-cycle cosine annealing scheduler. Hyperparameter tuning details, including model architectures and optimizer settings, are provided in Appendix 0.E, while Appendix 0.D describes the downsampling techniques applied in the dynamic experiments. All experiments are conducted over three independent trials with different random seeds to ensure robust evaluation, reporting the mean and standard deviation of Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as the primary evaluation metrics.
Table 1: Jammer localization for static scenarios, reporting mean MAE and RMSE over three trials with different seeds. Results are split by different sampling strategies, where circle (C), triangle (T), rectangle (R), and random (RD) denote geometric distributions of sampling points relative to the jammer.

### v2

@@ Abstract
Graph-based learning provides a powerful framework for modeling complex relational structures; however, its application within the domain of wireless security remains significantly underexplored. In this work, we introduce the first application of graph-based learning for jamming source localization, addressing the imminent threat of jamming attacks in wireless networks. Unlike geometric optimization techniques that struggle under environmental uncertainties and dense interference, we reformulate the localization as an inductive graph regression task. Our approach integrates structured node representations that encode local and global signal aggregation, ensuring spatial coherence and adaptive signal fusion. To enhance robustness, we incorporate an attention-based Graph Neural Network (GNN) that adaptively refines neighborhood influence and introduces a confidence-guided estimation mechanism that dynamically balances learned predictions with domain-informed priors. We evaluate our approach under complex Radio Frequency (RF) environments with various sampling densities, network topologies, jammer characteristics, and signal propagation conditions, conducting comprehensive ablation studies on graph construction, feature selection, and pooling strategies. Results demonstrate that our novel graph-based learning framework significantly outperforms established localization baselines, particularly in challenging scenarios with sparse and obfuscated signal information. Our code is available at https://github.com/tiiuae/gnn-jamming-source-localization.
@@ 1 Introduction
Graphs serve as a fundamental framework for representing complex relationships and interactions in real-world systems. Many highly successful machine learning applications are based on graph-based learning [23]. Although explored in numerous domains, the application of graph-based learning to wireless security remains underexplored. In particular, jamming source localization, a task essential for mitigating the threat of interference to the availability of wireless communication, presents a promising avenue for research. In this study, we investigate the application of a novel graph-based learning framework to perform an inductive graph-level regression task to predict the location of a jammer in complex RF environments.
@@
The ubiquitous and growing dependence on wireless networks for everyday connectivity and mission-critical operations introduces significant security vulnerabilities. As jamming attacks transmit intentional interference across communication channels, they lead to degradation and severing of wireless links [16]. The repercussions of such attacks are severe, leading to disruptions in essential services and operational hazards [28]. Although countermeasures such as frequency hopping have been explored, their effectiveness is limited against advanced jammers such as reactive, follow-on, and barrage that adaptively pursue target frequencies creating dense interference across multiple channels [17]. Unlike avoidance techniques that passively adjust the network to evade interference, jammer localization provides a direct mitigation strategy, wherein network administrators can deploy countermeasures such as physical neutralization or geofencing to restore network reliability regardless of the attack strategy [28].
@@
Classical jammer localization methods rely on geometric and optimization-based techniques to estimate the jammer’s position. However, they degrade in real-world RF environments due to noise and multipath effects [14]. Their dependence on idealized propagation models limits adaptability to stochastic RF dynamics. To overcome these limitations, we introduce a novel graph-based formulation of the jamming localization problem, leveraging attention-based GNNs to adaptively extract spatial and signal-related patterns from measurements. Additionally, we propose Confidence-guided Adaptive Global Estimation (CAGE), a confidence-guided estimation mechanism that dynamically balances GNN-based predictions with domain-informed priors, improving robustness in varying deployment conditions. Rather than treating localization as a geometric optimization problem, we redefine it as a graph regression task, where node features encode key RF and spatial characteristics, and the graph structure captures local and global signal dependencies. Our contributions are as follows:
@@
  • We present the first application of GNNs to jamming source localization in wireless networks by reformulating the problem as an inductive graph regression task.
@@ 2.1 Jamming Source Localization
Jamming source localization has been widely studied using range-free and range-based algorithms. Range-free algorithms perform localization using network topology-related properties, without relying on the physical characteristics of the incoming signals [11]. These methods are useful in infrastructure-limited environments, however, they suffer from degraded accuracy in sparse or unevenly distributed networks [25, 28]. In contrast, range-based methods utilize geometric optimization to estimate the distance to the source using the measurement of various physical properties such as Received Signal Strength Indicator (RSSI), Time of Arrival (ToA), and Angle-of-Arrival (AoA[11]. While these methods generally achieve higher accuracy, obtaining measurements such as ToA and AoA rely on specialized and calibrated hardware and are susceptible to errors where multipath effects introduce significant biases [29].
Many of the existing jammer localization approaches are validated under theoretical propagation models, such as free-space path loss, which fails to capture the complexities of real-world settings. Furthermore, many evaluations fail to include diversity in attack scenarios, sampling strategies, and long-range jamming effects [15]. To address these limitations, we evaluate our method on diverse network configurations, leveraging the Log Distance Path Loss (LDPL[6] model to account for realistic signal propagation conditions (See Appendix 0.D).
As our method leverages spatial and signal information exclusively, we focus on methods that similarly harness this information for localization. We benchmark our approach against established range-free and range-based localization techniques: Weighted Centroid Localization (WCL[25], Least Squares (LSQ[28], Path Loss (PL[13], Maximum Likelihood Estimation (MLE[13], and Multilateration (MLAT[30]. These baselines serve as reference points to assess the robustness and adaptability of our graph-based framework.
@@ 2.2 Graph Neural Networks for Localization
GNNs have demonstrated strong capabilities in regression tasks by capturing spatial dependencies within graph-structured data, such as in molecular property prediction and material science [12]. Beyond these domains, GNNs have also been applied to pose regression problems, including camera pose estimation [22] and human pose tracking [31], where they refine node and edge representations to improve motion prediction and spatial consistency. Their effectiveness in localization tasks has also gained attention, particularly in wireless and sensor network applications. GNNs have been applied to RF-based localization, including WiFi fingerprinting-based indoor localization [9], where they leverage RSSI signals to construct graphs and improve positioning accuracy. More recent works extend GNNs for network localization beyond Wi-Fi, addressing challenges such as dynamic network topologies and mobility-induced signal variation [6].
While these advancements highlight the growing potential of GNNs for network localization, their application to jamming source localization remains unexplored. This gap presents an opportunity to adapt GNNs to interference localization, addressing unique challenges under adversarial conditions.
@@ 3 Problem Definition
The problem addressed in this study is the localization of a wireless jammer, an adversarial interference source that disrupts communications by emitting intentional interference signals to degrade legitimate wireless communication.
3.0.1 Network Configuration
@@
where $s_{i}^{t}$ denotes a measurement taken at position $\mathbf{x}_{i}^{t}$ with an associated noise floor value $\eta_{i}^{t}$ in dBm at time $t$. Note that $\eta_{i}^{t}$, referred to as the noise floor, is commonly treated as jamming signal strength, where here it represents the combined effect of jammer interference and baseline environmental noise floor. Signal attenuation and propagation effects are modeled using the empirical LDPL model to simulate Non-line-of-sight (NLOS) conditions; refer to Appendix 0.D for details.
@@
3.0.2 Jammer Characteristics
The jammer is located at an unknown position $\mathbf{x}_{j}\in\mathcal{A}$ and emits interference signals that elevate the noise floor of nearby devices. The interference strength varies spatially due to distance attenuation and propagation effects such as shadowing and multipath fading. The sampled region, denoted as $\mathcal{R}\subseteq\mathcal{A}$, is the subset of $\mathcal{A}$ where devices collect noise floor measurements. Although the jammer’s position satisfies $\mathbf{x}_{j}\in\mathcal{A}$, it does not necessarily hold that $\mathbf{x}_{j}\in\mathcal{R}$, i.e., $\mathbf{x}_{j}\in\mathcal{A}\setminus\mathcal{R}$ is possible. We evaluate localization methods in scenarios where the jammer affects areas beyond the sampled region, testing their ability to infer positions outside direct measurement zones.
3.0.3 Objective
Given the set of samples collected by $N$ devices, denoted as $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\dots,\mathcal{S}_{N}\}$, the goal is to infer $\mathbf{x}_{j}$ based on the spatial distribution of measured noise floor levels. This problem is inherently challenging due to environmental noise and uncertainty affecting signal propagation, complex spatial correlations between noise floor levels and jammer interference, and the need to generalize beyond observed regions where no direct measurements are available.
@@ 4 Graph-Structured Learning for Jammer Localization
Given the set of collected measurements $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\dots,\mathcal{S}_{N}\}$, we represent the sampled signal space as a graph $G=(V,E)$, where nodes in $V$ correspond to individual measurement instances $s_{i}^{t}\in\mathcal{S}_{i}$, and edges $E$ define spatial relationships between them. Edges are established using K-Nearest Neighbors (KNN), where for each node $v_{i}$, an edge set is constructed as $E=\{(v_{i},v_{j})\mid v_{j}\in\mathcal{N}_{k}(v_{i})\}$, where $\mathcal{N}_{k}(v_{i})$ denotes the set of $k$ nearest neighbors of $v_{i}$ in the Euclidean space. To enforce the spatial attenuation principle of jamming signals, each edge $(v_{i},v_{j})$ is assigned a weight $w_{ij}$ that decays exponentially with the Euclidean distance as:
@@ 4.1 Learning Spatial Relations with Attention-Based Graphs
Graph-based learning enables the model to capture structured dependencies in the jammer interference field, allowing for improved generalization across spatial regions. However, effective learning in this setting requires handling two key challenges: (1) signal measurements are inherently noisy due to environmental conditions, leading to unreliable observations, and (2) the importance of neighboring nodes varies depending on both their spatial proximity and their reliability in capturing interference effects. Traditional GNN architectures, such as Graph Convolutional Networks [8], perform uniform neighborhood aggregation, which limits their ability to weigh the informativeness of spatially adjacent nodes. Our approach follows Graph Attention Network (GAT[24], incorporating an adaptive weighting mechanism that dynamically refines neighborhood influence, ensuring that message passing prioritizes nodes with reliable signal information while attenuating contributions from potentially misleading observations.
@@ 4.2 Supernode-Guided Adaptive Estimation
4.2.1 Global Context Encoding with Domain-Guided Priors
@@
Figure 1: Overview of the proposed jammer localization framework. A graph is constructed where nodes represent spatial and signal instances, and edges capture spatial relationships. The encoder processes the graph to learn spatial correlations. The final jammer position is estimated through an adaptive combination of the GNN prediction and WCL prior, controlled by a learned confidence weight.
4.2.2 Confidence-Guided Adaptive Position Estimation
@@
where $\boldsymbol{\alpha}\in\mathbb{R}^{5}$ is a five-dimensional confidence vector, with each $\alpha_{d}$ corresponding to one of the five output components in the normalized angular representation. The parameters $\mathbf{W}_{\alpha}\in\mathbb{R}^{5\times F^{\prime}}$ and $\mathbf{b}_{\alpha}\in\mathbb{R}^{5}$ are learnable, and sigmoid $\sigma(\cdot)$ ensures that $10<\alpha_{d}<$ 1. Finally, the predicted jammer position is computed as:
@@ 4.3 Training Strategy and Loss Function
4.3.1 Encouraging Independent Learning of the GNN Regressor
@@
Note that in the experimental evaluation, we refer to our proposed method as CAGE. For clarity in comparisons, Multilayer Perceptron (MLP), GCN, Principal Neighbourhood Aggregation (PNA) and GAT operate on graph $G$ and are trained with Equation (15) with the final estimate given by $\hat{\mathbf{x}}_{\text{GNN}}$, while CAGE is evaluated on the augmented graph $G^{\prime}$ and the adaptive confidence-weighted estimation, trained with Equation (14).
@@ 5 Experimental Evaluation
We evaluate CAGE across static and dynamic environments under the LDPL model. In the static setting (Section 5.1), fixed nodes with varying positions and densities are considered, with the jammer positioned either inside or outside the sampled region $\mathcal{R}$ to study spatial effects. In the dynamic setting (Section 5.2), a device moves in 3D space, approaching and encircling the jammer to assess localization accuracy across angles and distances. Ablation studies in Section 5.3 and Appendix 0.A evaluate the impact of node features, edge construction, graph augmentations, global pooling, and downsampling techniques. Details on data generation, node spatial arrangements, jammer characteristics, and signal propagation environments varied in our experiments are provided in Appendix 0.D, and an analysis of confidence weighting is presented in Section 5.3.
We compare CAGE against classical methods (WCL [28], LSQ [28], PL [13], MLAT [30], MLE [13], MLP) and graph-based learning methods (GCN [8], PNA [3], GAT [24]). As previously described, while MLP, GCN, PNA, and GAT operate on $G$, CAGE leverages augmented graph $G^{\prime}$ with the supernode, incorporating graph attention mechanisms alongside confidence-weighted estimation for adaptive localization. Models are trained using AdamW with a one-cycle cosine annealing scheduler. Hyperparameter tuning details, including model architectures and optimizer settings, are provided in Appendix 0.C. Appendix 5.3.4 describes the downsampling techniques applied in the dynamic experiments. All experiments are conducted over three independent trials with different random seeds to ensure robust evaluation, reporting the mean and standard deviation of Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as the primary evaluation metrics.
Table 1: RMSE in jammer localization for static scenarios, averaged over three trials with different seeds. Results are split by sampling geometry. MAE results are provided in Appendix 0.B Table 8.